OpenStudy (anonymous):

Find the difference quotient and simplify your answer. f(x) = x^2 - x + 1, f(5+h)-f(5)/h, h can not = 0

OpenStudy (anonymous):

\[ f(5+h) = (5+h)^2-(5+h)+1 \]

OpenStudy (anonymous):

Simplify that on its own first.

OpenStudy (anonymous):

Expand it.

OpenStudy (anonymous):

Expand it? I'm sorry I'm really horrible at math. So basically just multiply it out like... \[5f+hf=25+10h+h^2-(5+h)+1?\]

OpenStudy (anonymous):

yeah \[ 25+10h+h^2-5-h+1=h^2+9h+21 \]

OpenStudy (anonymous):

Ah, okay. That's it?

OpenStudy (anonymous):

Now what is \(f(5)\)?

OpenStudy (anonymous):

Wouldn't you replace f with 5? But in the equation you posted there's no f :/

OpenStudy (anonymous):

replace \(x\) with \(5\).

OpenStudy (anonymous):

oH x, that's what i meant, not f. but there's still no x in the equation you posted unless you mean replace the h? or are you telling me to do \[f(5)=x^2-x+1\] ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Replace x or the h? which did you say yes to? sorry :c

OpenStudy (anonymous):

Okay so far we already calculated \(f(5+h)\). Now we need to calculate \(f(5)\)

OpenStudy (anonymous):

\[ f(5)=(5)^2-(5)+1 \]

OpenStudy (anonymous):

Okay. So, \[f(5)=25-5+1 \] which is \[f(5) = 21\] ?

OpenStudy (anonymous):

Yes. Now we can put then both into : \[ \frac{f(5+h)-f(5)}{h} \]

OpenStudy (anonymous):

Ah so it would be \[(h^2+9h+21)-(21) / h ?\] and then... \[h^2+9h / h\] ?

OpenStudy (anonymous):

Yeah, now simplify.

OpenStudy (anonymous):

\[h+9\] ?

OpenStudy (anonymous):

In my book it says the answer should be \[3+h, h \neq 0\] how does it get to be 3?

OpenStudy (anonymous):

I'm not sure...

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=%5B%285%2Bh%29%5E2-%285%2Bh%29%2B1+-+%285%5E2-5%2B1%29%5D+%2Fh Our answer is correct. So their answer is wrong or our function is wrong.

OpenStudy (anonymous):

Ah, it must be a mistake because my friend had the same answer as well. Thank you so much!